کہا M sh 523 252 343 a 312 GIVEN: Vag = 608 Find: a) Reqab b)
Answer #1

M50 4. 52 312 Vab = 60V Reqab S Reqab 50 is in series with so or is in series with 40 6+4 = 100m, 5+5=10 wlos — — — -b ton 32Now lon is in halallel with lon of loxlo lo tlo 10o = 52 20 - ses § 52 - LM TAA 5 r _ 312 is in series with 312 5731=361 362Regal = 180 || 360 188 36 - 648 _ 18 +36 : 54 Reqab = 120 - ___ Given Vab= lov Consider the final - circuit 361 Gov 1893 13tr850 łz6r 3lr + Ven : 5rx isa uso e sx izse (6.360 is the series - combination of 312 R ands) . user= 5x1.667 usa : 8.335 v nConsider the previous circuit - b lour slou Goy tail - use=you (:51 is the - parallel combination of your andlor) C - Vor=8.3